EVEN POWERS OF DIVISORS AND ELLIPTIC ZETA VALUES

GIOVANNI FELDER ALEXANDER VARCHENKO

Abstract. We introduce and study elliptic zeta values, a two-parameter deformation of the values of Riemann’s zeta function at positive integers. They are essentially Taylor coefficients of the logarithm of the elliptic gamma function, and share the SL(3, ℤ) modular properties of this function. Elliptic zeta values at even integers are related to Eisenstein series and thus to sums of odd powers of divisors. The elliptic zeta values at odd integers can be expressed in terms of generating series of sums of even powers of divisors.